# Mixing, segregation, and collapse transitions of interacting copolymer rings

### v1

@@ 5.2 SE-M phase boundary
If $\beta_{c}=0$ then the model reduces to two polygons with a mutual nearest neighbour attraction between vertices in different polygons. In this case the model goes through a segregated-mixed phase transition at $\beta_{m}^{*}\approx 0.31$ [JOTW22]. The average number of mutual contacts per unit length, $\langle k_{m}\rangle/2n$, is shown as a function of $\beta_{m}$ (for $\beta_{c}=0$ and $0.28$) in figure 7 for $n=48,100,148,200,300$ and $400$.
@@
Figure 8: (a) The locations of the crossings of the variance per unit length curves $Var(k_{m})/2n$ as a function of $n=400$ with the corresponding curves for $m=48,148,200,300$, (b) As in the left panel but for the curves $d_{cm}$ against $\beta_{m}$. The three values of $\beta_{c}$ are $0.0,0.20$ and $0.28$ (see legend). In both panels data have been extrapolated against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ is the estimate of $\beta_{m}^{*}$ (see reference [JOTW22]) for the $\beta_{c}=0$ case.
@@
Figure 10: Left panel: The locations of crossings $\beta_{m}^{*}(Var(k_{m})^{cr})$ between the curve $Var(k_{m})/2n$ for $n=400$ (see the middle row of figure 9 and $Var(k_{m})/2n$ for $n\in\{48,100,148,200,300\}$ for values of $\beta_{c}<\beta_{c}^{*}$ (below the $\theta$-point; these are the empty symbols) with those above $\beta_{c}^{*}$ (solid symbols). The locations are plotted as a function of $\sqrt{n}$. Notice that the data below the $\theta$-point tends to line up and extrapolate as $n\to\infty$ to the marked solid vertical bar at $1/\sqrt{n}=0$ (this is the estimate for $\beta_{m}^{*}$ in reference [JOTW22] for the case $\beta_{c}=0$. In contrast, the data above the $\theta$-point extrapolate to larger values on the vertical axis. Right panel: This is similar to the left panel, but now for the curves $d_{cm}$ in the borrow row of figure 9. The values of $\beta_{c}$ are reported in the legend.
The onset of a thermodynamic phase transition between the SE- and M-phases is also suggested by the $\beta_{m}$ dependence of $Var(k_{m})/2n$ in the curves in the middle row of panels in figure 7. Notice the onset of peaks whose heights increase with $n$. Moreover, the curves cross one another in a narrow region of the mutual interaction parameter $\beta_{m}$. This is consistent with the theorem that, for $\beta_{c}<\infty$, and $\beta_{m}<\beta_{m}^{*}$, $\lim_{n\to\infty}Var(k_{m})/2n=0$; see section 3 and theorem 1. That result shows the presence of an asymmetric phase transition when $\beta_{c}<\infty$ somewhat similar to the adsorption transition of a self-avoiding walk anchored to an impenetrable attractive wall [HTW82]. In these cases it is known that the intersections of the curves $Var(k_{m})/2n$ plotted against $\beta_{m}$ provide a faithful finite-size estimate of the location of the phase boundary [buks_book_2015].
@@
To determine the approximate location of the critical point $\beta_{m}^{*}$ the crossings between the curves between $Var(k_{m})/2n$ and $d_{cm}$ were determined and plotted against $1/\sqrt{n}$ in figure 8 (assuming the mean field value $1/2$ for the finite size crossover exponent $\phi$ and plotting the intersections between the variance for $n=400$ and for $n=48,100,148,200$ and $300$). Extrapolating these data points indicates a limiting value close to $0.31$ (the thick segment at $1/\sqrt{n}=0$ marks the estimate of the expanded-mixed phase transition $\beta_{m}^{*}=0.31\pm 0.01$ reported in reference [JOTW22]). A similar extrapolation is reported for the location of the crossings of the curve $d_{cm}(n=400)$ with the corresponding curves obtained for $n=48,100,148,200$ and $300$ in the last row of panels in figure 8. The limiting estimate shown by the thick segment at $1/\sqrt{n}=0$ is again the expanded-mixed phase transition point $\beta_{m}^{*}=0.31\pm 0.01$ reported in reference [JOTW22]. Similarly to the SE-SC phase boundary, we note that: (i) the $n\to\infty$ extrapolation of the points of both measures is very close to $\beta_{m}^{*}$, and (ii) the SE-M phase transition appears to be independent of the strength $\beta_{c}$ of the self attraction between vertices in the polygons, at least when $0\leq\beta_{c}<\beta_{c}^{*}$.
Figure 12: (a) Comparison of the locations of the maxima of the variance curves $Var(k_{c})/2n$ for various values of $\beta_{m}$. These data extrapolate to $\beta_{m}^{*}$ for data below the M-SE transition (this corresponds to the empty symbols), but to larger values above the M-SE transition (solid symbols) the M-SE transition. (b) Similar to (a), not now for the separation between the centre-of-mass data $d_{cm}$ plotted for various values of $\beta_{m}$. The values of $\beta_{m}$ are reported explicitly in the legend. In both panels data have been plotted as a function of $1/\sqrt{n}$. The vertical bars at $1/\sqrt{n}=0$ refers to the estimates of $\beta_{m}^{*}$ provided in reference [JOTW22] for the $\beta_{m}=0$ case.
@@ 5.4 The M-SC phase boundary
Distance between the two centre of mass:
@@
Estimates of the location of the M-SC phase boundary, based on the crossings of the $Var(k_{c})$ curves and the $d_{cm}$ curves are plotted in figure 12. These points are again plotted against $1/\sqrt{n}$, and the vertical bar on the $y$-axis is the confidence interval of estimate of $\beta_{m}^{*}$ in reference [JOTW22] when $\beta_{c}=0$.
@@ 6 Topological entanglement
Linking between the $A$- and $B$-blocks is computed via the two-variable Alexander polynomial $\Delta(t,s)$ [torres1953alexander] while knot detection is based on the computation of the Dowker code [dowker1983classification].
Since their computation would be prohibitively costly when the number of crossings $n_{c}$ after a planar projection is very large, we first simplify the geometrical entanglement of the system while keeping its topology unaltered [van1990knot, van2002probability]. This is achieved by shrinking stochastically the polygons with the BFACF algorithm, simulated at a very small step fugacity $K$ [1981bergrandom, de1983new, marco2014knotted]. The BFACF moves have the property of preserving the topology and, if $K$ is sufficiently small, the system will often eventually shrink to its minimal length compatible with its topology. Even if the minimal length is not always reached, the shrunk configuration, once projected onto a plane, displays, most of the time, a sufficiently small number of crossings (well below $50$) to dramatically reduce the time required to compute the topological invariants [van1990knot, millett2013identifying].
@@ 6.1 Fixing $\beta_{m}$ and varying $\beta_{c}$
Knotting of the component polygons should increase when the SE-SC phase boundary is crossed, and this is seen in the bottom row of panels in figure 13. The knotting probability increases from a small value to significant values in the SC phase, consistent with previous studies showing low knotting probabilities of expanded polygons in the cubic lattice, which increase rapidly as the $\theta$-point is crossed into a phase of collapsed polygons [tesi1994knotting, mansfield2007development].
@@ 7 Discussion
In section 4, we give a brief description of the Monte Carlo method we use to investigate the phases in the phase diagram. We performed extensive simulations using a Multiple Markov Chain implementation [geyer1995annealing, tesi1996monte, tesi1996interacting] of the Metropolis Algorithm [metropolis1953monte] on lattice polygons using pivot elementary moves [madras1990monte] and local Verdier-Stockmayer style moves [verdier1962monte].

### v2

@@ 5.2 SE-M phase boundary
If $\beta_{c}=0$ then the model reduces to two polygons with a mutual nearest neighbour attraction between vertices in different polygons. In this case the model goes through a segregated-mixed phase transition at $\beta_{m}^{*}\approx 0.31$ [12]. The average number of mutual contacts per unit length, $\langle k_{m}\rangle/2n$, is shown as a function of $\beta_{m}$ (for $\beta_{c}=0$ and $0.28$) in figure 7 for $n=48,100,148,200,300$ and $400$.
@@
Figure 8: (a) The locations of the crossings of the variance per unit length curves $Var(k_{m})/2n$ as a function of $n=400$ with the corresponding curves for $m=48,148,200,300$, (b) As in the left panel but for the curves $d_{cm}$ against $\beta_{m}$. The three values of $\beta_{c}$ are $0.0,0.20$ and $0.28$ (see legend). In both panels data have been extrapolated against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ is the estimate of $\beta_{m}^{*}$ (see reference [12]) for the $\beta_{c}=0$ case.
@@
Figure 10: Left panel: The locations of crossings $\beta_{m}^{*}(Var(k_{m})^{cr})$ between the curve $Var(k_{m})/2n$ for $n=400$ (see the middle row of figure 9) and $Var(k_{m})/2n$ for $n\in\{48,100,148,200,300\}$ for values of $\beta_{c}<\beta_{c}^{*}$ (below the $\theta$-point; these are the empty symbols) with those above $\beta_{c}^{*}$ (solid symbols). The locations are plotted as a function of $1/\sqrt{n}$. Notice that the data below the $\theta$-point tends to line up and extrapolate as $n\to\infty$ to the marked solid vertical bar at $1/\sqrt{n}=0$ (this is the estimate for $\beta_{m}^{*}$ in reference [12] for the case $\beta_{c}=0$. In contrast, the data above the $\theta$-point extrapolate to larger values on the vertical axis. Right panel: This is similar to the left panel, but now for the curves $d_{cm}$ in the middle row of figure 9. The values of $\beta_{c}$ are reported in the legend.
The onset of a thermodynamic phase transition between the SE- and M-phases is also suggested by the $\beta_{m}$ dependence of $Var(k_{m})/2n$ in the curves in the middle row of panels in figure 7. Notice the onset of peaks whose heights increase with $n$. Moreover, the curves cross one another in a narrow region of the mutual interaction parameter $\beta_{m}$. This is consistent with the theorem that, for $\beta_{c}<\infty$, and $\beta_{m}<\beta_{m}^{*}$, $\lim_{n\to\infty}Var(k_{m})/2n=0$; see section 3 and theorem 1. That result shows the presence of an asymmetric phase transition when $\beta_{c}<\infty$ somewhat similar to the adsorption transition of a self-avoiding walk anchored to an impenetrable attractive wall [10]. In these cases it is known that the intersections of the curves $Var(k_{m})/2n$ plotted against $\beta_{m}$ provide a faithful finite-size estimate of the location of the phase boundary [16].
@@
To determine the approximate location of the critical point $\beta_{m}^{*}$ the crossings between the curves between $Var(k_{m})/2n$ and $d_{cm}$ were determined and plotted against $1/\sqrt{n}$ in figure 8 (assuming the mean field value $1/2$ for the finite size crossover exponent $\phi$ and plotting the intersections between the variance for $n=400$ and for $n=48,100,148,200$ and $300$). Extrapolating these data points indicates a limiting value close to $0.31$ (the thick segment at $1/\sqrt{n}=0$ marks the estimate of the expanded-mixed phase transition $\beta_{m}^{*}=0.31\pm 0.01$ reported in reference [12]). A similar extrapolation is reported for the location of the crossings of the curve $d_{cm}(n=400)$ with the corresponding curves obtained for $n=48,100,148,200$ and $300$ in the last row of panels in figure 8. The limiting estimate shown by the thick segment at $1/\sqrt{n}=0$ is again the expanded-mixed phase transition point $\beta_{m}^{*}=0.31\pm 0.01$ reported in reference [12]. Similarly to the SE-SC phase boundary, we note that: (i) the $n\to\infty$ extrapolation of the points of both measures is very close to $\beta_{m}^{*}$, and (ii) the SE-M phase transition appears to be independent of the strength $\beta_{c}$ of the self attraction between vertices in the polygons, at least when $0\leq\beta_{c}<\beta_{c}^{*}$.
Figure 12: (a) Comparison of the locations of the maxima of the variance curves $Var(k_{c})/2n$ for various values of $\beta_{m}$. These data extrapolate to $\beta_{m}^{*}$ for data below the M-SE transition (this corresponds to the empty symbols), but to larger values above the M-SE transition (solid symbols) the M-SE transition. (b) Similar to (a), not now for the separation between the centre-of-mass data $d_{cm}$ plotted for various values of $\beta_{m}$. The values of $\beta_{m}$ are reported explicitly in the legend. In both panels data have been plotted as a function of $1/\sqrt{n}$. The vertical bars at $1/\sqrt{n}=0$ refers to the estimates of $\beta_{m}^{*}$ provided in reference [12] for the $\beta_{m}=0$ case.
@@ 5.4 The M-SC phase boundary
Distance between the two centres of mass:
@@
Estimates of the location of the M-SC phase boundary, based on the crossings of the $Var(k_{c})$ curves and the $d_{cm}$ curves are plotted in figure 12. These points are again plotted against $1/\sqrt{n}$, and the vertical bar on the $y$-axis is the confidence interval of estimate of $\beta_{m}^{*}$ in reference [12] when $\beta_{c}=0$.
@@ 6 Topological entanglement
Linking between the $A$- and $B$-blocks is computed via the two-variable Alexander polynomial $\Delta(t,s)$ [31] while knot detection is based on the computation of the Dowker code [7].
Since their computation would be prohibitively costly when the number of crossings $n_{c}$ after a planar projection is very large, we first simplify the geometrical entanglement of the system while keeping its topology unaltered [13, 15]. This is achieved by shrinking stochastically the polygons with the BFACF algorithm, simulated at a very small step fugacity $K$ [4, 1, 2]. The BFACF moves have the property of preserving the topology and, if $K$ is sufficiently small, the system will often eventually shrink to its minimal length compatible with its topology. Even if the minimal length is not always reached, the shrunk configuration, once projected onto a plane, displays, most of the time, a sufficiently small number of crossings (well below $50$) to dramatically reduce the time required to compute the topological invariants [13, 22].
@@ 6.1 Fixing $\beta_{m}$ and varying $\beta_{c}$
Knotting of the component polygons should increase when the SE-SC phase boundary is crossed, and this is seen in the bottom row of panels in figure 13. The knotting probability increases from a small value to significant values in the SC phase, consistent with previous studies showing low knotting probabilities of expanded polygons in the cubic lattice, which increase rapidly as the $\theta$-point is crossed into a phase of collapsed polygons [30, 20].
@@ 7 Discussion
In section 4, we give a brief description of the Monte Carlo method we use to investigate the phases in the phase diagram. We performed extensive simulations using a Multiple Markov Chain implementation [8, 29, 28] of the Metropolis Algorithm [21] on lattice polygons using pivot elementary moves [19] and local Verdier-Stockmayer style moves [33].